# A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies

### v2

@@ Abstract
Understanding how social norms vary across cultures can help us build culturally aligned NLP systems. We propose a culture-agnostic approach to norm discovery, using moral emotions, shame and pride, to identify examples of normative expectations and extract corresponding social norms. We present the first cross-cultural self-conscious emotions dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over 10K extracted social norms. We validate our dataset using native speakers and demonstrate how our dataset reveals variations in social norms that align with the cultural dichotomy observed in these nations -- e.g., Bollywood movies emphasize shame due to deviation from social roles, and express pride in family honor, while Hollywood shames poverty and incompetence, and takes pride in ethical behavior. Notably, females are shamed more across both cultures and both cultures shame women for violating similar normative expectations.
A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies
@@ 1 Introduction
Normative expectations, [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al. (1990) use injunctive norms. We use the terms social norms and normative expectations interchangeably. commonly called ‘‘social norms’’, refer to people’s beliefs about what others approve of or expect people to do. These norms play a significant role in shaping and regulating social behavior and can reveal underlying moral and ethical values.
Recent research on social alignment of LLMs has either crowd-sourced social norms from QnA style posts on social media platforms or prompted pre-trained LLMs to extract them Jiang et al. (2021); Nahian et al. (2020); Fung et al. (2022). However, a majority of these studies are conducted in English and capture Western social situations. Furthermore, LLMs such as GPT-4 OpenAI (2023) are predominantly trained on data from Western cultures, Havaldar et al. (2023b), and thus, the generated "social norms" tend to be Anglocentric.
@@
Okay, I am going.
$\Rightarrow$ Incestuous relationship evokes shame
@@
Table 1: Excerpts of dialogues with target search words.
@@
Self-conscious emotions, or moral emotions, are emotional reactions to a violation or an observance of normative expectations in society. In turn, these emotions reflect social sanctions and approvals. Unlike basic emotions such as joy and anger, self-conscious emotions such as shame and pride are evoked by self-reflection and self-evaluation. Examining situations evoking self-conscious emotions can thus reveal unspoken social norms specific to a culture (See Table 1). While the function (i.e., social regulation) of self-conscious emotions is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022), their appraisals and expressions vary Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022).
@@
In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, (b) propose a culture-agnostic approach for norm discovery and release a dataset of $10K$ normative expectations and, (c) examine cross-cultural variations in social norms. We study India and the United States of America to investigate the utility of moral emotions for cross-culture norm discovery. India is predominantly a collectivist society where one’s sense of self is interwoven with community beliefs, whereas the USA is predominantly an individualist society that values competency and autonomy Triandis (1989, 1988). The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their moral and ethical values.
Relatedly, gender differences in social behaviors emerge from societal expectations and the process of socialization. For instance, women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We examine how shame, an indicator of negative evaluation, is attributed to women across cultures. We hypothesize that more shame is attributed to women in Indian society due to its patriarchal nature compared to American society.
Movies Total Tokens #shame #pride #control
Bollywood 2738 22.62M 5409 2999 8303
Hollywood 2697 20.78M 1221 2805 4385
5435 43.3M 6583 5804 12688
Table 2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
2 Background
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. SOCIAL-CHEM-101 Forbes et al. (2020), SCRUPLES Lourie et al. (2021), SOCIAL BIAS FRAMES Sap et al. (2019), MORAL INTEGRITY CORPUS Ziems et al. (2022), and VALUE PRISM Sorensen et al. (2023) are a few datasets developed to teach socially aligned interactions to LLMs. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics. Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al. (2023b); Atari et al. (2023). Capturing diversity in morality is thus a research question of great significance Talat et al. (2021).
In this paper, we examine the cross-cultural variations in self-conscious emotions and use these emotions to extract culture-specific normative expectations. We validate our findings by evaluating their alignment with known cultural tendencies of Indian and American societies. Additionally, we examine gender-wise attribution of shame and norm distribution to contextualize the potential impact of inculcating automatically extracted social values into machines.
3 Dataset
Culture-specific people’s conversations are needed to study moral emotions and to mine underlying normative expectations. Social media QnA posts are primarily recollections or ’what-if’ situations that do not reveal fully how these emotions are expressed in natural conversations, for instance between a boss and an employee or among family members at a Christmas table.
@@ 3 Dataset
Movies are a rich source of real-time social interactions illustrating power dynamics, negotiations, and self-conscious emotions in action; to study how and why these emotions are expressed. We collected English subtitles for 5,435 movies belonging to Bollywood and Hollywood that were released post-1990 by automatically crawling websites that host or link movie subtitles (See Table 2 for movie distribution). The choice of English language for Bollywood movies was made based on the availability of subtitles. The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
3.1 Data Proprocessing
We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context. One author manually checked the entire dataset of shame and pride-related dialogues to correct spelling mistakes in subtitles, remove unrecognizable characters, and identify phrases having pride and shame-related keywords but not indicating social sanction (or approval).
For instance, the phrases like "what a shame" or "it’s a shame" in Hollywood movies do not reflect social disapproval, and therefore such dialogues were excluded. Below is an example of such dialogue.
@@ 3.1 Data Proprocessing
Similarly, the phrase "proudly presents" was used in the context of introducing an event or an artist in Hollywood movies. All dialogues with the phrase "proudly presents" were thus removed for further analysis.
We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Additionally, we also formed a control set of dialogues unrelated to shame and pride for both movie industries (See Table 2 for dialogue distribution).
Table 3: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with shame for Bollywood and Hollywood after Benjamini-Hochberg p-value correction. Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood. See Table A6 and A8 for the complete set of correlations.
@@ 4 Approach
Fig. 1 illustrates our study design comprising two key approaches (a) a vocabulary approach to measure cross-cultural variations in linguistic expressions of shame/pride and (b) prompting LLM approach to extract latent normative expectations and examine how they differ between India and the U.S.A.
@@ 4.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological properties of words such as "self-focus" (e.g., I, my, mine) vs. "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we extracted the normalized distribution of psycho-social categories from the dialogues (shame+control vs pride+control for both movie industries) and examined their correlation with shame and pride. Please note that search keywords (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
4.2 Pretrained LLM: GPT-4
Taking into account pre-trained LLMs’ tendency to favor an Anglocentric understanding of text Havaldar et al. (2023b), we designed the prompts to elicit the "reason behind self-conscious emotions" rather than extracting "social norm" to overcome the cultural bias. The prompts were designed to seek the following:
@@ 4.2 Pretrained LLM: GPT-4
The first question orients the LLM to focus on the subject, i.e., person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social norm in the culture. Since GPT-4 exhibits performance at par with human annotators Pan et al. (2023); Törnberg (2023), we designed this task using GPT-4 chat in a two-shot setting. The temperature was set to 00. The prompts with two-shot samples are provided in Tables A4 and A5.
Themes in Social Norms
To understand differences in situations that evoke shame and pride in Indian and American societies, we performed agglomerative clustering after embedding extracted norms using SBERT embeddings Reimers and Gurevych (2019). For this purpose, we merged shame-related norms from Bollywood and Hollywood and pride-related norms from Bollywood and Hollywood and removed duplicates. Social norms extracted from GPT-4 chat are short phrases (See Table 5) and are devoid of culture-specific style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
Table 4: Pearson r with $95\%$ confidence interval for top LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction. See Tables A7 and A9 for all significantly correlated categories.
@@ 5 Results
5.1 Cross-cultural variations in expressions of shame and pride
Shame
The selected top LIWC categories significantly correlated with shame in Bollywood and Hollywood is provided in Table 3. Although shame is associated with negative emotions, power, and morality in both industries (See Table A6 and A8 for all significant correlations), significant cultural nuances exist in its manifestation. In Hollywood, shame is I-focused with remorse (e.g., sadness, negative emotions, anxiety), whereas in Bollywood, it is You-focused with more social references (e.g., you, your, he, her). Shame is past-focused in Hollywood, whereas it is present-focused in Bollywood movies. Anxiety is unique to Hollywood-shame whereas anger is more strongly associated with shame in Bollywood. Categories such as female, sexual, and swearing are exclusively correlated with shame in Bollywood indicating a conservative society whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
Pride-
related interactions are family focused and self-centered (e.g., I am proud of __) in Hollywood movies whereas it is achievement/power focused and "We"-centered in Bollywood movies (See Table 4). Similarly, pride expressions related to power, morals, and prosocial behavior are more common in Bollywood dialogues, whereas pride expressions related to individuals’ growth (i.e., male/female, work) are more frequent in Hollywood dialogues. Both industries associate pride with ethnicity.
5.2 Cross-cultural variations in Social norms
Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing the self-conscious emotions (shame and pride)" and the reason behind self-conscious emotions in a randomly sampled set of $100$ dialogues from Hollywood and $100$ dialogues from Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
Gender was incorrect for eight samples in the Bollywood set and for five samples in the Hollywood set. For Bollywood, GPT-4 chat predicted a gender for 15 (3 predicted as Female, rest as male) samples where gender was not evident, whereas, for 10 (3 Female) such samples in Hollywood. One such example is provided below:
e.g. We’re proud of Bristol’s decision to have her baby… and even prouder to become grandparents? We’re not proud that… our teenage daughter is pregnant, Maria. And I don’t want to send a message that teen pregnancy… is something to be proud of. I want that line out. Of course, Governor.
Here, the gender is predicted as "Male" for Governor. GPT-4 chat returns the most likely gender indicating bias however, it is not always incorrect. It is thus safe to assume that prediction error for gender lies between 8-23% for Bollywood and 5-15% for Hollywood.
For Bollywood, 11 predicted reasons behind self-conscious emotion were incorrect whereas for Hollywood, 2 such samples were incorrect. The reason was not explicit for another nine samples in Bollywood set and one sample in Hollywood. Overall, the prediction error lies between 11-20% for Bollywood whereas it is 2-3% for Hollywood.

### v3

# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

@@ Abstract
Social emotions such as shame and pride reflect social sanctions or approvals in society. In this paper, we examine how expressions of shame and pride vary across cultures and harness them to extract unspoken normative expectations across cultures. We introduce the first cross-cultural shame/pride emotions movie dialogue dataset, obtained from  5.4K Bollywood and Hollywood movies, along with over $10K$ implicit social norms. Our study reveals variations in expressions of social emotions and social norms that align with known cultural tendencies observed in the United States and India -- e.g., Hollywood movies express shame predominantly toward self whereas Bollywood movies express shame predominantly toward others. Similarly, Bollywood shames non-conformity in gender roles, and takes pride in collective identity, while Hollywood shames lack of accountability, and takes pride in ethical behavior. More importantly, women face more prejudice across cultures and are sanctioned for similar social norms.
Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice
@@ 1 Introduction
Normative expectations [^1] Bicchieri (2005) uses the term normative expectations, Lahti and Weinstein (2005) use moral norms, and many social psychologists Cialdini et al. (1990) use injunctive norms. We use the terms social norms and normative expectations interchangeably., a category of social norms that refer to others’ beliefs about how we ought to behave Bicchieri (2005). It reveals attitudes and values that influence social actions across cultures. For instance, hesitancy toward seeking mental help could exist because others expect an individual not to seek help or because the individual believes that others do not seek help. The former reveals the attitude of society, whereas the latter indicates an individual’ lack of awareness. Capturing diversity in cultural beliefs is paramount to building socially aligned AI models for diverse users Talat et al. (2021); Atari et al. (2023), for inducing collective behavior change (e.g., overcoming stigma against mental health disorders) and several other social applications.
@@
$\Rightarrow$ Incestuous relationship evoke shame.
@@
Table 1: Excerpts of dialogues expressing explicit social emotions shame and pride indicating actions and social approval.
Prior works crowd-sourced social norms from Q&A style social media posts (such as Reddit) Jiang et al. (2021); Nahian et al. (2020) or prompted large language models (LLMs) Fung et al. (2022) without distinguishing between normative and empirical expectations. However, social media platforms [^2] Twitter Statistic: https://blog.hootsuite.com/twitter-statistics/, [^3] Reddit Statistic: https://www.enterpriseappstoday.com/stats/reddit-statistics.html are skewed toward Western demographics (i.e., young white male) and the reliance on a small cohort of annotators typically located in Western countries further enforces Anglocentric beliefs when collecting socially acceptable behaviors Ziems et al. (2022). LLMs with sophisticated capabilities such as GPT-4 are increasingly used to extract social norms; however, Havaldar et al. (2023a) illustrated their lack of cross-cultural understanding.
@@
The overarching question we ask in this paper is, how and what social behaviors are sanctioned or encouraged across cultures. Shame and pride are social emotions experienced across cultures on violating or adhering to normative expectations in society Tracy and Robins (2004). Notably, the function of shame (i.e., social regulation) is consistent across cultures Goetz and Keltner (2007); Fessler (2007); Schaumberg and Skowronek (2022). Examining situations evoking shame and pride can thus reveal unspoken social norms specific to a culture (See Table 1).
@@
In this paper, we (a) examine cross-cultural variations in linguistic expressions of shame and pride, and (b) propose a cross-cultural social emotion-based approach for norm discovery. The appraisals and expressions of shame and pride vary across cultures and is a research question of great significance to understand society and self-regulation Wong and Tsai (2007); Lewis et al. (2010); Wegge et al. (2022). We chose two top-grossing movie industries, i.e., Hollywood and Bollywood, to analyze how social emotions and norms differ across cultures. Hollywood primarily depicts social situations from the USA, an individualist society that values competency and autonomy Triandis (1989, 1988) whereas Bollywood depicts India, a collectivist society where one’s sense of self is interwoven with community beliefs. The cultural dichotomy between India and the USA (i.e., collectivism vs. individualism) thus presents a rich ground for understanding variations in their beliefs and values.
We release a large cross-cultural dataset of social emotions-related dialogues with $10K$ social norms. Using a blend of psychology-informed language analysis and LLM-powered approach to mine implicit information, we demonstrate cross-cultural variations in how and why these emotions are expressed. In parallel, we examine how social sanctions are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We hypothesize that women are more subjected to social sanctions than men across cultures.
2 Movies Subtitles Corpora
Movies provide life-like depictions of social situations to study how and why these emotions are expressed. They are influenced by and/or, known to influence real lives Adkins and Castle (2014); Kubrak (2020). In contrast, social media Q&A posts are (a) skewed to social situations specific to tech-savvy young demographic groups, and (b) are second-hand reports that do not reveal fully the social dynamics of how these emotions are expressed in natural conversations, for instance, between a boss and an employee or among family members at a Christmas table.
@@ 2 Movies Subtitles Corpora
For our study, we collected English subtitles for 5,435 Hollywood and Bollywood movies that were released post-1990 by auto-crawling websites that host or link movie subtitles (See Table A2 for data distribution). The year of release for movies was verified by either parsing subtitle file names having a release year or checking Wikipedia entries. The year mapping was performed to ensure a similar period for collected movies.
2.1 Extracting Shame and Pride related Expressions
The expressions of shame and pride could be implicit (as indicated by lowered shoulders, avoiding eye contact, etc.), and explicit (See Table 1). Detecting implicit social emotions in textual discourse is an innately complex task requiring socio-cultural knowledge that may not be evident to outsiders or third-party annotators. Below is one such example:
‘‘Actor 1: Liza, Alice! Come away from that window this minute. I think it’s disgusting, standing there where they can see you.
Actors 2,3 (Liza, Alice): I notice that you peek out often enough when you think no one’s looking.
Actor 1: I never!
Actors 2,3 (Liza, Alice): I saw you! So did I.
Here, it is difficult to determine if shame is evoked and/or experienced by actors in this discourse. A third-party annotation will reveal their beliefs and attitudes, which could introduce bias in our analysis. Adopting a high precision-low recall strategy, we focus on explicit mentions of shame and pride to learn about occurrences of shame and pride as determined by the characters. We extracted dialogues having the word shame or pride or their variations (See Appendix A1), along with the previous and the next two lines for situational context (See Table 1). In the case of short dialogues, which could be the case due to monosyllabic responses in spoken conversations, an extra previous and next line was appended to context.
Manual Screening One author manually checked the entire dataset of shame and pride-related dialogues to filter out dialogues unrelated to social sanctions/approvals e.g., phrases such as "what a shame", "it’s a shame", "proudly presents". Below is an example of such dialogue.
@@ 2.1 Extracting Shame and Pride related Expressions
Similarly, the phrase "proudly presents" is used while introducing an event or an artist in Hollywood movies.
Final Dataset: We created four sets of dialogues: (a) shame-related dialogues in Bollywood, (b) shame-related dialogues in Hollywood, (c) pride-related dialogues in Bollywood and (d) pride-related dialogues in Hollywood. Cross-cultural differences also exist in speaking styles across cultures and we thus formed a control set of dialogues unrelated to shame and pride for both movie industries to remove the variations in language markers owing to culture-specific speaking styles (See Table A2 for dialogue distribution).
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Moral (wrong, excuse, decent, honest, duty) 0.110 $2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.161
$1^{st}$ person sing. pronouns (i, me, i’m, my, i’ll) 0.102 Social References (you, your, he, her, him ) 0.148
Sadness (crying, cry, sob, lonely, sad) 0.096 Moral (wrong, innocent, duty, decent, excuse) 0.102
$2^{nd}$ person pronouns (you, your, you’re, yourself, you’ve) 0.080 Female ( her, she, girl, she’s, mom) 0.060
Past focus (was, did, were, been, didn’t ) 0.069 Family (son, married, uncle, dad, mom) 0.055
All or none (no, all, nothing, never, yes) 0.055 Sexual (chaste, lust, sex, sexy, pimp) 0.051
Family (son, dad, baby, mom, mama) 0.054 Conflict (kill, killed, accusing, killing, cruel) 0.024
Social References (you, your, we, he, you’re) 0.050 Swear words (hell, bloody, idiot, damn, ass) 0.024
Anxiety (fear, afraid, worry, terrified, scared) 0.042 Anger (mad, angry, hate, cruel, argue) 0.022
Illness (sick, pain, pains, flu, sickly) 0.038 Present focus (is, are, don’t, i’m, aren’t) 0.021
Table 2: Pearson r for LIWC categories significantly correlated ($p<0.05$) with shame for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast that is, Self-centered shame vs Other-centered shame; Past focus vs Present focus in Hollywood vs Bollywood. See Table A7 and A9 for the complete set of correlations.
@@ 3 Approach
Fig. 1 illustrates our study design comprising a blend of two key approaches (a) a psychosocial vocabulary approach to measure how expressions of shame and pride differ and (b) prompting Large Language Model approach to extract reasons behind them and how they differ between India and the U.S.A.
@@ 3.1 Vocabulary Approach: LIWC
Linguistic Inquiry of Word Count (LIWC) Boyd et al. (2022) is a corpus analysis tool widely used in psychology to identify psychological categories of words such as "self-focus" (e.g., I, my, mine) and "other-focus" (e.g., you, yours, theirs) which are of great significance for examining social behaviors such as self-regulation and conformity.
To understand cross-cultural linguistic variations in the manifestation of shame and pride, we computed the normalized distribution of psycho-social categories from the dialogues and examined their correlation with shame and pride compared to the control set. The search keywords used for building social emotions corpus (See Table A1) were removed from the LIWC dictionary to prevent overestimation of shame- (e.g., negative emotion) and pride-related categories (e.g., achievement).
3.2 Pretrained LLM: GPT-4 Chat
LLMs such as GPT-4 exhibit superlative English language understanding and are more reliable than human annotators on platforms such as MTurk Pan et al. (2023); Törnberg (2023). We used GPT-4 chat in two-shot setting ( See Tables A5 and A6 for prompts) to seek answers to the following questions:
@@ 3.2 Pretrained LLM: GPT-4 Chat
Since there are at least two characters in a discourse, the first question orients the LLM to focus on the person experiencing the social sanction or approval and then identify their gender. The output for the second question serves as the implicit social norm in the culture. Asking for "reason" leading to the expression of social emotion encourages LLMs to retrieve norm from provided context, mitigating potential Anglocentric tendencies.
Thematic Analysis of Social Norms
To capture overarching themes in reasons behind expressing shame and pride in Indian and American societies, we performed agglomerative clustering after embedding unique shame and pride-related social norms using SBERT embeddings Reimers and Gurevych (2019). Unlike movie dialogues, social norms extracted from GPT-4 chat are short phrases (See Table 4) and are devoid of culture-specific language style markers (i.e., Indian English vs American English) thus, no control for culture is needed.
Hollywood: LIWC Category (Top-5 words) $r$ Bollywood: LIWC Category (Top-5 words) $r$
Social References (you, your, we, he, you’re) 0.131 Drive (we, our, us, sir, work) 0.122
Family (son, dad, baby, mom, mama) 0.121 Achieve (work, better, win, best, try) 0.098
$1^{st}$ person sing. pronouns (i, i’m, me, my, i’ll) 0.095 Power (sir, own, respect, kill, power) 0.091
$2^{nd}$ person pronouns (you, your, you’re, you’ve, yourself) 0.091 Social References (you, your, he, we, our) 0.085
Male References (he, his, him, man, son) 0.084 Moral (wrong, duty, brave, arrogant, useless) 0.075
Drive (we, our, us, we’re, dad) 0.076 Reward (win, won, glory, success, successful) 0.071
Ethnicity (American, Irish, Chinese, German, Christian) 0.052 Political (nation, army, sultan, president, dynasty) 0.064
Reward (win, won, winner, successful, earned) 0.048 $1^{st}$ person pl. pronouns (we, our, us, we’ll, let’s) 0.060
Achieve (work, better, best, trying, try) 0.045 Family (son, papa, married, dad, uncle) 0.059
Power (sir, own, war, strong, mighty) 0.044 Ethnicity (Indian, Indians, British, Hindi, caste) 0.054
Table 3: Pearson r for LIWC categories significantly correlated ($p<0.05$) with pride for Hollywood and Bollywood after Benjamini-Hochberg p-value correction (See Appendix for confidence intervals and p values). Note the contrast, Achievement-related pride in Bollywood, Self-centered pride vs We-centered pride, Family in Hollywood vs Family in Bollywood See Tables A8 and A10 for all significantly correlated categories.
@@ 4 Results
4.1 How do expressions of shame/pride vary across cultures?
Shame is associated with negative emotions, power, and morality in both movie industries (See Table 2), however, significant cultural nuances exist in its manifestation. In Hollywood, shame is more I-focused with remorse (e.g., sadness, anxiety) and past-focused language, whereas in Bollywood, it is more You-focused with frequent social references and present-focused language. In Bollywood, psycho-social categories associated with shame reveal a larger social context (note the contrast in family i.e., nuclear family in Hollywood vs extended family in Bollywood) and social situations (i.e., anger, swear, conflict). Categories female and sexual are exclusively correlated with shame in Bollywood potentially indicating an honor system in collectivist societies Caffaro et al. (2014) whereas it is illness that is uniquely correlated with shame in Hollywood indicating social sanctions around incapability.
A similar dichotomy is observed for Pride related interactions (See Table 3). Pride in Hollywood interactions is more centered on family (note, family in Bollywood is male-centric and comprises extended family (such as uncle)) and self-centered (e.g., I am proud of __) whereas it is achievement/power focused and "We"-centered in Bollywood movies (with othering (us vs they)- See Appendix). Male references are more strongly associated with pride than females in Hollywood whereas the female category has an insignificant ($p>0.05$) correlation with pride in Bollywood. Ethnicity-related pride is expressed in both movie industries but political category is exclusively correlated to Bollywood. Pride is more positive (see POSEMO in Table 3) in Hollywood whereas in Bollywood, it reflects honor (e.g., protecting family’s pride, bringing pride to family).